OpenStudy (anonymous):

Factor AND Solve. 10x - 50 = 0

OpenStudy (solomonzelman):

Factor out of 10, and divide both sides by 10. Add 5 to both sides, and tell me what x equals to.

OpenStudy (anonymous):

So I divide 10 and 50 by 10?

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf 10x - 50 = 0 }\) \(\LARGE\color{blue}{ \bf 10(x - 5) = 0 }\) Divide each side by 10. Divide 10(x-5) by 10, and you will get x-5, and divide 0 by 10, and you will get 0 So... \(\LARGE\color{blue}{ \bf x - 5 = 0 }\) Now, can you do the last step?

OpenStudy (anonymous):

5.

OpenStudy (anonymous):

Factor AND Solve. a² + 9a = 0

OpenStudy (solomonzelman):

well, factor out of a. \(\LARGE\color{purple}{ \bf a^{2} +9a=0 }\) factor out of a... \(\LARGE\color{purple}{ \bf a(a +9)=0 }\) divide both sides by a... \(\LARGE\color{purple}{ \bf a +9=0 }\) do the last step for me please. \(\LARGE\color{purple}{ \bf a =? }\)

OpenStudy (anonymous):

I'm so confused.

OpenStudy (anonymous):

Is it -9?